WAC Notes March 7 2020

From PREX Wiki
Jump to: navigation, search
Documentation
HOW TOs for shift crew
Expert Tools
All Expert Contacts



PREX Main << Weekly Analysis Coordinator << WAC Notes CREX << Instructions

WAC Notes March 6 2020 << >> WAC Notes March 8 2020

March 7 2020

Daily Meeting BlueJeans


Production Run List - Slug 157 - IHWP OUT - Flip-LEFT - Production

Slug started at approx 03:00 on March 7

Production Run List - Slug 158 - IHWP IN - Flip-LEFT - Production

Slug started at approx 21:30 on March 7

  • Good - Run 6991
    • Current steps up from 120 uA to 130 uA to 140 UA during run
    • BCM worries: Looks good. Pretty trippy, so BCM DD mean value jumps around. Concern is from digitals BCM DD disagreeing with analogs (and being a 2 sigma outlier) > So the digitals probably need to be checked on for this run. Specifically BCM DG US disagrees the most w.r.t. the rest of them and it has 5 extra ppm of noise for some reason
  • Good - Run 6992
    • 140 uA
  • Good - Run 6994
    • 140 uA
  • Good - Run 6995
    • 140 uA
  • Good - Run 6996
    • 140 uA
  • Good - Run 6997
    • 140 uA
  • Good - Run 6998
    • 140 uA
  • Good - Run 6999
    • 140 uA
  • Good - Run 7000
    • 140 uA
  • Good - Run 7001
    • 140 uA
  • Good - Run 7002
    • 140 uA
  • Good - Run 7003
    • 145 uA
  • Good - Run 7004
    • 143 uA

Non-Production Runs and Bad Runs

  • Run 6979
    • Junked for having no beam
  • Runs 6984-6986
    • Junked for injector work, beam recovery
  • Run 6993
    • Down for a period of time, re-established on C-1%.

Issues

  • Noticed temperature in LHRS hut was rising, called in techs to fix during day shift.
  • FFB became unstable as slug 157 wore on.

Updates